OpenStudy (anonymous):

What is the surface area of the cone? A. 115π cm^2 B. 120π cm^2 C. 90π cm^2 D. 65π cm^2

OpenStudy (ybarrap):

Do you know the formula?

OpenStudy (anonymous):

nope :/

OpenStudy (jdoe0001):

total surface area for a cone => \(\bf \pi r \sqrt{r^2+h^2}+\pi r^2\)

OpenStudy (ybarrap):

where l is the side length of your cone

OpenStudy (ybarrap):

Do you know what "r" is in your figure?

OpenStudy (anonymous):

the height is 12

OpenStudy (anonymous):

radius is 5?

OpenStudy (ybarrap):

yes, and yes

OpenStudy (ybarrap):

formula produces the answer like magic

OpenStudy (anonymous):

65π+ 25π= 90π?

OpenStudy (ybarrap):

yep

OpenStudy (anonymous):

thank you :)

OpenStudy (ybarrap):

np